A brass bar, 15 cm in length, is welded to an aluminum bar, 10 cm in length. Both have a cross section of 3 cm on a side. If the free end of the brass bar is at 50 degrees Celcius and the the free end of the aluminum bar is at 25 degrees Celcius, what is the steady-state temperature at the junction of the 2 bars? *




1
Expert's answer
2021-06-24T16:14:45-0400

Solution.

"l_b=0.15m;"

"l_a=0.1m;"

"d=0.03m;"

"t_b=50^oC;"

"t_a=25^oC;"

"\rho_b=8500kg/m^3;"

"\rho_a=2700kg/m^3;"

"c_b=380J\/kg^oC;"

"c_a=920J\/kg^oC;"

"m_bc_b\Delta T_b=m_ac_a\Delta T_a;"

"m_b=\rho_bS_bl_b; m_a=\rho_aS_al_a;"

"\rho_bl_bc_b\Delta T_b=\rho_al_ac_a\Delta T_a;"

"\dfrac{\Delta T_b}{\Delta T_a}=\dfrac{\rho_al_ac_a}{\rho_bl_bc_b};"

"\dfrac{\Delta T_b}{\Delta T_a}=\dfrac{2700\sdot0.1\sdot920}{8500\sdot0.15\sdot380}=0.5;"

"\dfrac{t_b-t_f}{t_f-t_a}=0.5\implies t_f=\dfrac{t_b+0.5t_a}{1.5};"

"t_f=\dfrac{50+12.5}{1.5}=41.7^oC;"

Answer: "t_f=41.7^oC."